Solve for x and y

√ x + y =11
x+√y=7
1
Expert's answer
2011-08-11T10:06:33-0400
√ x + y =11
x+√y=7

x = (11 - y)2
y = (7 - x)2

x = 121 - 2y +y2
y = 49 - 2x + x2

y = 49 - 2(121 - 2y +y2)+ (121 - 2y +y2)2

-y4 + 4y3 - 244y2 + 481y = 14448
This equation has complex solutions.